Foam Products, Inc., makes foam seat cushions for the automotive and aerospace industries. The company’s activity-based costing system has four activity cost pools, which are listed below along with their activity measures and activity rates:

Activity Cost Pool Activity Measure Activity Rate
Supporting direct labor Number of direct labor-hours $ 14 per direct labor-hour
Batch processing Number of batches $ 88 per batch
Order processing Number of orders $ 282 per order
Customer service Number of customers $ 2,637 per customer

The company just completed a single order from Interstate Trucking for 2,000 custom seat cushions. The order was produced in four batches. Each seat cushion required 0.3 direct labor-hours. The selling price was $141.20 per unit, the direct materials cost was $107 per unit, and the direct labor cost was $14.20 per unit. This was Interstate Trucking’s only order during the year.

Required:

Calculate the customer margin on sales to Interstate Trucking for the year.

Interstate Trucking
Customer Margin—ABC Analysis
Costs:
Customer margin
Answer #1
Interstate Trucking
Customer Margin—ABC Analysis
Sales 282400 =2000*141.20
Costs:
Direct materials 214000 =2000*107
Direct labor 8520 =2000*0.3*14.20
Supporting direct labor 8400 =2000*0.3*14
Batch processing 352 =88*4
Order processing 282
Customer service overhead 2637
234191
Customer margin 48209